# A Gentle Wakeup Call: Symmetry Breaking with Less Collision Cost

### v1

@@ Abstract
The wakeup problem addresses the fundamental challenge of symmetry breaking. There are $n$ devices sharing a time-slotted multiple access channel. In any fixed slot, if a single device sends a packet, it succeeds; however, if two or more devices send, then there is a collision and none of the corresponding packets succeed. For the static version of wakeup, all packets are initially active (i.e., can send and listen on the channel); for the dynamic version, the packets become active at arbitrary times. In both versions, the goal is to successfully send a single packet.
Prior results on wakeup have largely focused on the number of slots until the first success; that is, the latency. However, in many modern systems, collisions introduce significant delay, an aspect that current wakeup algorithms do not address. For instance, while existing results for static wakeup have polylogarithmic-in-$n$ latency, they can incur additional latency that is linear in the cost of a collision $\mathcal{C}$. Thus, the total latency is large and dominated by the contributions from collisions.
Here, we design and analyze a randomized wakeup algorithm, Aim-High. For sufficiently large $\mathcal{C}$ and with bounded error, Aim-High has latency and expected collision cost that is nearly $O(\sqrt{\mathcal{C}})$ for both the static and dynamic versions. Otherwise, the latency and expected collision cost are $O({\texttt{poly}({\log n})})$ for the static setting, and $O(n\,{\texttt{poly}({\log n})})$ for the dynamic setting. We also establish lower bounds that complement these results.
Keywords:
Wakeup, symmetry breaking, multiple access channel, collision cost
@@ 1 Introduction
In the wakeup problem, devices are classified as either active (awake) or inactive (asleep). Each active device has a packet to send on a shared channel, while inactive participants are limited to listening. The first successful transmission acts as a “wakeup call” where all sleeping participants are woken up, and the problem is solved [2].
Traditionally, prior work on this problem has focused on the time until the first success, which is typically referred to as the latency. Minimizing this metric is challenging, since in any unit of time, the shared channel can only accommodate a single sender; if two or more devices send simultaneously, the result is a collision where none of the corresponding packets are successfully sent. Moreover, sending on the channel is performed in a distributed fashion; that is, a priori there is no central scheduler or coordinator.
Under the standard cost model for wakeup, each collision wastes a single slot. However, in many settings, a collision can cost us much more than a single slot. For instance, in mobile systems, a failed communication due to a collision may result in the sender having to wait until the intended receiver is again within transmission range [26]. For WiFi networks, a collision is known to consume time proportional to the packet size [1]. Therefore, collisions may significantly impact the performance of modern systems.
In this work, we address both latency and collision cost in solving the wakeup problem. As we will discuss later in Section 3.1, while existing algorithms and natural approaches yield strong latency guarantees, they often result in high collision costs. Therefore, optimizing for these two metrics requires a new approach.
@@ 1.1 Model and Notation
Communication. Time is divided into disjoint slots, each of which can accommodate a packet transmission. Communication occurs on a multiple access channel, which is defined as follows. In any slot, a packet can either attempt to transmit or listen to the channel. If no packet is transmitted in a given slot, it is empty. If only one packet is transmitted, the packet succeeds; we often refer to this as a success. However, if multiple packets transmit simultaneously, they all fail—this is a collision—and they may try again later. There is no ability to distinguish between an empty slot and a collision; that is, there is no collision detection. There are no identifiers for the packets. Finally, packets must be sent in a distributed manner; that is, there is no scheduler or central authority.
Packet Activation. We consider two settings. For the static setting, in the first slot, all $n$ packets are active on the multiple access channel and begin executing an algorithm. By implication, all packets agree on the index of the current slot; this feature is typically referred to as a global clock [2].
In the dynamic setting, all $n$ packets become active over time. To capture the worst case, we consider these activation times to be scheduled by an adversary, and we often refer to packets being activated or injected by the adversary. Packets may not agree on the index of the current slot, and thus there is no global clock.
Metrics. We care about latency, which refers to the time between the first packet injection and the first success. Additionally, each collision incurs a known cost of $\mathcal{C}$ $\geq 4$. [^1] The value $4$ is chosen to simplify the analysis. We also note that only $\mathcal{C}$ $\omega(1)$ is interesting; otherwise, an optimal algorithm (such as sawtooth backoff [16, 13, 2]) for the standard model will be asymptotically optimal. To capture the worst-case performance, $\mathcal{C}$ is chosen by the adversary with knowledge of $n$, subject to $\mathcal{C}=O({\texttt{poly}({n})})$, where the degree is of the polynomial is unknown. Note that this reveals no information about $n$, whose value is unknown to the packets a priori.
Given any wakeup algorithm, the pertinent costs are: (i) the collision cost, which is the number of collisions multiplied by $\mathcal{C}$, and (ii) the latency. Although both metrics represent time-based costs and we aim to minimize their maximum, we report them separately for clarity and to facilitate our lower-bound results.
[…]
For a random variable $X$, we say that w.h.p. in $x$ the expected value $E[X]\leq v$ if, when tallying the events in the expected value calculation, the total probability of those events where $X>v$ is $O(1/{\texttt{poly}({x})})$, where the constant degree of the polynomial is can be made arbitrarily large.
Finally, our results are parameterized by $\epsilon$, which is a small positive constant $\ll 1$ set by our algorithm prior to execution. For terms of the form $\log^{k+(k^{\prime}/\epsilon)}n$, where $k,k^{\prime}$ are positive constants at least $1$, we report this as $\log^{\Theta(1/\epsilon)}n$ for ease of presentation.
@@ 1.2 Our Results
Our algorithm, Aim-High (AH), applies to the static and dynamic wakeup problems. [^2] We explain the name in Section 3.1 We first address the static setting, which is closely tied to the classic problem of leader election (see Section 2). Here, we define $\eta$ $\max\{n,\mathcal{C}\}$, which is used when bounding the error of our upper bounds derived in Section 4.1.
@@
W.h.p. in $\eta$, AH solves the static wakeup problem. The following associated costs hold w.h.p. in $\eta$:
@@
    If  $\mathcal{C}=\Omega(\log^{\Theta(1/\epsilon)}n)$, then the latency and the expected collision cost are each $O\left(\mathcal{C}^{\frac{1}{2}+2\epsilon}\right)$.
@@
    Else, the latency and expected collision cost are $O(\log^{\Theta(1/\epsilon)}n)$.
By way of comparison, existing approaches to wakeup incur a collision cost that is $\Omega(\mathcal{C})$ (see Section 3.1). In contrast, Theorem 1.1 implies that if $\mathcal{C}$ is “large”, AH offers performance that is significantly sublinear in $\mathcal{C}$. Otherwise, the latency and the expected collision cost are each polylogarithmic-in-$n$, which echoes prior wakeup results without collision detection (see Section 2).
Many upper bounds in the static setting are achieved by “fair” algorithms. A fair algorithm is defined by a sequence $p_{1}$, $p_{2}$, $\dots$, where $p_{i}$ denotes the sending probability used by every packet in the $i$-th slot. In Section 4.2, we give the following lower bound for this class of algorithms (to which AH belongs) having a w.h.p. bound on the expected latency.
@@
Let $\alpha$ and $\beta$ be constants such that $1/(2\epsilon)\geq\alpha\geq 0$ and $\beta\geq 0$. Any fair algorithm that w.h.p. in $\eta$ solves the static wakeup problem with $O(\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n)$ latency has w.h.p. an expected collision cost of $\Omega(\mathcal{C}^{(1/2)-k\epsilon}/\log^{\beta}n)$.
This theorem, along with noting that the bound in (2) of Theorem 1.1 is within a polylogarithmic-in-$n$ factor of the optimal latency and expected collision cost, yields the following result.
Corollary 1
When $\mathcal{C}=\Omega(\log^{\Theta(1/\epsilon)}n)$, AH is within $O(\mathcal{C}^{4\epsilon})$-factor of the optimal expected collision cost for fair algorithms, and within a polylogarithmic-in-$n$ factor of the optimal latency and expected collision cost otherwise.
In Section 4.3, we note that while AH upper bounds the latency and expected collision cost with w.h.p. in $\eta$, if this guarantee fails, then these quantities may be unbounded. We describe revisions to AH that yield the same bounds for expected latency and expected collision cost that hold with probability $1$.
Next, we analyze the performance of AH in the dynamic setting, and we give the following upper-bound results. Here, our w.h.p. guarantees are in terms of $\mathcal{C}$.
Theorem 1.3
W.h.p. in $\mathcal{C}$, AH solves the dynamic wakeup problem. The following associated costs hold w.h.p. in $\mathcal{C}$:
  1. 1.
    If  $\mathcal{C}=\Omega(n^{2-\delta}\log^{\Theta(1/\epsilon)}n)$ for a small constant $\delta>0$, then the latency and the expected collision cost are each $O\left(\mathcal{C}^{\frac{1}{2}+\epsilon}\log n\right)$,
  2. 2.
    Else, the latency is $\tilde{O}(n(\log^{\Theta(1/\epsilon)}n))$ and the expected collision cost is $\tilde{O}(n(\log^{\Theta(1/\epsilon)}n))$.
Again, if $\mathcal{C}$ is large, then we achieve performance that is sublinear in $\mathcal{C}$. We give a lower bound for a class of algorithms we call “batch-fair" (to which AH belongs) that is analogous to fair algorithms, but natural in the dynamic setting.
Theorem 1.4
When $\mathcal{C}=\Omega(n^{2-\delta})$ for a small constant $\delta>0$, AH is within an $O(\mathcal{C}^{2\epsilon}\log^{2}n)$-factor of the optimal expected collision cost for batch-fair algorithms, and within an $O(\log^{\Theta(1/\epsilon)}n)$-factor of the optimal latency otherwise.
@@ 2 Related Work
The wakeup problem, has been extensively studied for decades. Here, we summarize closely-related results.
Static Wakeup. In the literature, static wakeup, also known as leader election, addresses the fundamental problem of symmetry breaking. Early work by Willard [25] gives an upper bound of $\lg\lg n+O(1)$, and a matching lower bound for fair algorithms with collision detection (CD).
@@
Given access to $b>1$ multiple channels with CD, Fineman et al. [12] give an algorithm with $O((\log n/\log b)+(\log\log n)(\log\log\log n))$ latency whp for any number of active devices. In the SINR model [22], Fineman et al. [11] demonstrate $O(\log n+\log R)$ latency w.h.p., where $R$ is the ratio of longest to shortest link, implying $O(\log n)$ latency as typically $R=O({\texttt{poly}({n})})$; they also provide a general $\Omega(\log n)$ lower bound. Chang et al [7] focus on energy efficiency under different models of channel feedback. Gilbert et al. [14] analyze the impact of system-size predictions, deriving lower and upper bounds.
Dynamic Wakeup. The static and dynamic wakeup problems differ significantly. Specifically, for dynamic wakeup without CD, Jurdzinski and Stachowiak [19] show an $\Omega(\log^{2}n)$ lower bound, which contrasts with the $\Omega(\log n)$ lower bound [21] and matching upper bound [3] in the static setting.
Without CD, Jurdziński and Stachowiak [18] explore the implications of global and local clocks, known or unknown $n$, packet labels, both with and without CD. Most relevant to our work is an algorithm that succeeds with probabilty at least $1-\varepsilon$ and has $O(n/\log n)$ latency when there is no knowledge of $n$, no global clock, no CD, and no labels; the probability of failure for the corresponding algorithm is parameterized
Under the SINR model without a global clock, assuming devices know a polynomial upper bound on $n$, Jurdzinski and Stachowiak [19] give a Monte Carlo algorithm with $O(\log^{2}n/\log\log n)$ latency w.h.p., and a Las Vegas algorithm with expected $O(\log^{2}n/\log\log n)$ latency, compared to a deterministic $O(\log^{2}n)$ latency [17]. Without CD and no global clock, Chlebus et al. [9] explore the impact of multiple $b>1$ channels, where they give a randomized algorithm with error $\varepsilon$ and latency $O(n^{1/b}\ln(1/\varepsilon))$.
Finally, we highlight prior work that addresses the cost of collisions in the context of the closely-related problem of contention resolution (see [8, 2]), where the goal is to have all $n$ packets successfully sent . Here, Biswas et al. [5] solve the static setting of contention resolution with latency and expected collision cost $\tilde{O}(n\sqrt{\mathcal{C}})$; however, they do not address the dynamic setting. Furthermore, Biswas et al. [5] assume CD, which their algorithm relies on to adjust the sending probability; this additional channel feedback is unavailable in our model. Another related result is by Anderton et al. [1] who show that collisions significantly degrade the performance of popular backoff algorithms for contention resolution.
3 Algorithm Specification
@@ 5 $w_{\mbox{\tiny cur}}\leftarrow w_{0}$
8 for slot $i=1$ to $d\sqrt{\mathcal{C}}\,\ell$ do
9
10   Send packet with probability $1/w_{cur}$
@@ 14    $w_{\mbox{\tiny cur}}\leftarrow w_{\mbox{\tiny cur}}/2$
15
16// Doubling Phase
17 while True do
@@ 18 $w_{\mbox{\tiny cur}}\leftarrow 2w_{\mbox{\tiny cur}}$
19 for slot $i=1$ to $d\ell$ do
20
21   Send packet with probability $1/w_{cur}$
@@ 23     Terminate
The pseudocode for AH is given above, and each packet executes this algorithm upon being activated. Under AH, each packet begins with an initial window size of $w_{0}$ $2^{\mathcal{C}^{\epsilon}}$; recall that $\epsilon$ $\ll 1$ is a small positive constant that influences our bounds on cost.
For the static setting, $\ell$ $\ln(w_{\mbox{\tiny cur}})$, where $w_{\mbox{\tiny cur}}$ is the current window size. For the dynamic setting, $\ell=\ln(\mathcal{C})$.
Initially, the current window size $w_{\mbox{\tiny cur}}$ is set to $w_{0}$, and this is updated throughout the execution of two phases, as we describe below and illustrated in Figure 1.
@@
Doubling phase. At the beginning of this phase, the window size doubles such that now $w_{\mbox{\tiny cur}}=4$ (Line 1). Then, each packet sends with probability $\Theta(1/w_{\mbox{\tiny cur}})$ in each slot of $d\ell$ slots (Line 1 and 1); we refer to these slots as a small sample. If a success occurs, the execution terminates; otherwise, this process continues.
@@ 3.1 Context and Overview
In this section, we start by considering how prior approaches fare in a model when there is a collision cost. Then, we discuss our design choices regarding AH, along with an overview of our analysis.
@@
In the static setting, a natural approach for breaking symmetry is to “backoff”, where each packet sends with probability $1/2^{i}$ in slot $i\geq 0$. At $i=\lg(n)$, the probability of a success in a slot becomes constant. This approach is a common ingredient in algorithm design; indeed, AH makes use of it. Another example of its use in solving wakeup is by Jurdziński and Stachowiak [18], who achieve $O(n/\log n)$ latency with bounded error $\varepsilon$ in a similar model to ours, but without a notion of collision cost. However, in this prior result, each active packet starts with a constant sending probability, which inflicts $\Omega(\mathcal{C})$ cost.
So, we can observe that while backoff may be an important part of any solution, a key obstacle is how to set the initial sending probability suitably low. We do not know $n$ a priori, and so it seems that no matter what initial sending probability we choose—even if it depends on the known value of $\mathcal{C}$—the adversary may choose a sufficiently large value of $n$ such that collisions are unavoidable. What can we do?
Figure 1: Illustration of the phases for AH. The top portion depicts the halving phase, where the initial window size is $w_{0}=2^{\mathcal{C}^{\epsilon}}$. The bottom portion depicts the doubling phase, starting from a window of size $4$, where the degree to which packets contend for channel access decreases as the window grows, leading to a likely success.
@@
Setting an Initial Window Size. The initial sending probability needs to be set carefully. Too high and we incur many (potentially expensive) collisions as we have seen above.
Consider a window of size $w^{*}=\Theta(n\sqrt{\mathcal{C}})$; we refer to this as a good window. Informally, once packets reach this good window, the chance of a success in any slot of the window is at least $\binom{n}{1}(1/w^{*})\approx 1/\sqrt{\mathcal{C}}$. Therefore, over $\Theta(\sqrt{\mathcal{C}})$ slots, we expect a success. At the same time, the probability of a collision is approximately $\binom{n}{2}(1/w^{*})^{2}\approx 1/\mathcal{C}$, and thus the expected collision cost per slot is $1$, and $O(\sqrt{\mathcal{C}})$ over $\Theta(\sqrt{\mathcal{C}})$ slots. This observation dictates the size of our large samples (recall Section 3).
In designing AH, we hope that $w^{*}\leq w_{0}$, since in this case, the above informal analysis yields good expected collision cost (and, as we will show, good latency). However, since the adversary sets $\mathcal{C}$ with knowledge of $n$, this is by no means guaranteed; the adversary can always make collisions overwhelmingly likely. However, in this case, our consolation prize is that collisions should be cheap. To elaborate, we choose to "aim high" with the initial window size; that is, setting $w_{0}=2^{\mathcal{C}^{\epsilon}}$. This choice sets up our analysis of the following two cases for the static setting:
@@
    Go Left For Good Window. If $n\sqrt{\mathcal{C}}\leq w_{0}$, then the good window exists at or to the left of the initial window; that is, the good-window size is less than or equal to the initial-window size. Here, the probability of a collision is “not too high”. Thus, so long as we adjust the sending probability upwards in a conservative manner, we can reduce expected collision cost; this is the purpose of the halving phase. Of course, there is a tension with latency, since this adjustment cannot be too conservative, otherwise the time until the first success will be too large.
@@
    Go Right For Good Window Conversely, if $n\sqrt{\mathcal{C}}>w_{0}$, then the good window exists to the right of the initial window. Here, the collision probability may be “too high”. However, a key realization here is that the collision cost is cheap, since we will argue that $n>2^{\mathcal{C}^{\epsilon}}/\sqrt{\mathcal{C}}$ implies that $\mathcal{C}=O(\log^{1/\epsilon}n)$. Thus, we can tolerate these collisions as we increase the window size in the doubling phase.
While the above cases address the particulars of our approach to the static setting, analogous intuition holds for the dynamic version (see Section 5.1 for discussion).
Samples. As discussed above, we "aim high" with the initial window and execute the the halving phase; here, we are hoping to succeed via samples of size $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$. However, if the adversary chooses to set $\mathcal{C}$ such that $n\sqrt{\mathcal{C}}>w_{0}$, good contention will not be discovered by the halving phase, and our ambitious setting of $w_{0}$ does not yield a success in the halving phase.
In this case, however, collisions are now provably cheap, and we no longer care about finding a good window; we simply want a quick success, even if we rack up collisions in the process. Thus, the doubling phase—which is classic backoff—uses smaller samples of size $d\log(w_{\mbox{\tiny cur}})$, and once a window of size $\Theta(n)$ is reached, a success is likely.
Finally, we note that in the static setting, the start and end points of each sample are the same for all packets. However, for the dynamic setting, this may not hold given that the adversary controls when packets are activated; that is, the samples of packets can be unsynchronized, and this poses an additional challenge.
Parameter $\epsilon$. Throughout, we treat $\epsilon$ as a small constant for the purposes of our analysis. Yet, it is worth pointing out how it interacts with our metrics of interest: latency and expected collision cost. The former is improved by making $\epsilon$ small, since this shortens the halving phase. So, if aiming high pays off, then a smaller $\epsilon$ can reduce latency. However, we pay for this in terms of the latter metric; specifically, when $n\sqrt{\mathcal{C}}>w_{0}$ and the algorithm looks to succeed in the doubling phase, we have $\mathcal{C}=O(\log^{1/\epsilon}n)$. Thus, degree of the polynomial increases with the reciprocal of $\epsilon$, which means our expected collision cost increases.
4 The Static Setting
In this section, we analyze the performance of AH for the static wakeup problem. Throughout our analysis, due to space constraints, some proofs have been moved to the appendix.
4.1 Upper Bound
Throughout, we assume that $n$ is sufficiently large. We make use of the following well-known facts.

### v2

# A Gentle Wakeup Call Symmetry Breaking with Less Collision Cost

@@ Abstract
The wakeup problem addresses the fundamental challenge of symmetry breaking. Initially, $n$ devices share a time-slotted multiple access channel, which models wireless communication. A transmission succeeds if exactly one device sends in a slot; if two or more transmit, a collision occurs and none succeed. The goal is to achieve a single successful transmission efficiently.
Prior work on wakeup primarily analyzes latency—the number of slots until the first success. However, in many modern systems, each collision incurs a nontrivial delay, $\mathcal{C}$, which prior analyses neglect. Consequently, although existing algorithms achieve polylogarithmic-in-$n$ latency, they still suffer a delay of $\Omega(\mathcal{C})$ due to collisions.
Here, we design and analyze a randomized wakeup algorithm, Aim-High. When $\mathcal{C}$ is sufficiently large with respect to $n$, Aim-High has expected latency and expected total cost of collisions that are nearly $O(\sqrt{\mathcal{C}})$; otherwise, both quantities are $O({\texttt{poly}({\log n})})$. Finally, for a well-studied class of algorithms, we establish a trade-off between latency and expected total cost of collisions.
@@ 1 Introduction
In the wakeup problem, devices are either active (awake) or inactive (asleep). Each active device has a packet to send on a shared channel, while inactive devices are limited to listening. The first successful transmission serves as a “wakeup call”, awakening all inactive devices and resolving the problem [2].
Prior work on this problem has focused on the time until the first success, which is typically referred to as the latency. Minimizing this metric is challenging, since in any unit of time, the shared channel can only accommodate a single sender; if two or more devices send simultaneously, the result is a collision where none of the corresponding packets are successfully sent. Sending on the channel is performed in a distributed fashion; that is, a priori there is no central scheduler or coordinator.
Under the standard cost model for wakeup, each collision wastes a single slot. However, in many settings, a collision can consume much more than a single slot. Therefore, collisions may significantly impact the performance of modern systems.
In this work, we address both latency and the cost of collisions required to solve the wakeup problem. While existing algorithms give strong latency guarantees, they result in a large number of collisions (see Section 3.1), and addressing these two metrics in tandem requires new ideas.
Why Should We Care About Collisions? The delay incurred by a collision can vary across different systems. As an example, consider WiFi (IEEE 802.11) networks; here, both sending a packet and the delay from a collision consumes time, say $\mathcal{C}$, proportional to the packet size [1]. In this setting, wakeup incurs an unavoidable delay of $\mathcal{C}$, since a single success is required. Yet, reducing the delay caused by collisions is still important. If this delay is, say, $X$, then throughput is $\mathcal{C}/(\mathcal{C}+X)$. So, if $X=\Theta(\mathcal{C})$, then throughput is sub-optimal by a constant factor. However, if say $X=\Theta(\sqrt{\mathcal{C}})$, then throughput is $\mathcal{C}/(\mathcal{C}+\Theta(\sqrt{\mathcal{C}}))\geq 1-\Theta(1/\sqrt{\mathcal{C}})$; that is, throughput approaches the optimal value of $1$ as $\mathcal{C}$ grows large. [^1] Empty slots will also degrade throughput, but in WiFi, these slots are negligible compared to the time of sending a packet or a collision. This illustrates how minimizing collisions improves performance.
More generally, over long-range links, where the propagation delays are large, a collision means devices can wait a long round-trip time before realizing the transmission has failed. Also, in intermittently-connected mobile wireless networks [21], failed communication due to a collision may result in the sender having to wait until the intended receiver is again within transmission range. Finally, in environments where there is some fluctuation in the delay (e.g., arising from fluctuating channel conditions or periods of heavy network usage), $\mathcal{C}$ can be interpreted as an upper bound on this delay that holds under most operating conditions.
@@ 1.1 Model and Notation
Communication. Time is divided into disjoint slots, each of which can accommodate a packet transmission. Communication occurs on a multiple access channel, which is defined as follows. In any slot, a packet can either attempt to transmit or listen to the channel. If no packet is transmitted in a given slot, it is empty. If only one packet is transmitted, the packet succeeds; we often refer to this as a success. However, if multiple packets transmit simultaneously, they all fail—this is a collision—and they may try again later. There is no scheduler or central authority a priori.
No Collision Detection. There is no mechanism to distinguish between an empty slot and a collision—that is, no collision detection (no-CD). In many practical systems, collision detection (CD) is unavailable or unreliable. For example, in WiFi networks, received signal strength may offer a weak form of CD, but it is error-prone. Consequently, many prior results address the challenging no-CD setting (see the survey [2] and related discussion in Section 2).
We emphasize that CD and the cost of collisions are orthogonal concerns: collisions incur a performance penalty regardless of whether they are detectable. For example, in WiFi, even if devices cannot sense an ongoing collision, the channel remains unavailable for its duration.
Metrics. We care about latency, which refers to the time between the first packet injection and the first success. Additionally, each collision incurs a known delay of $\mathcal{C}$ $\geq 4$. [^2] The value $4$ is chosen to simplify the analysis. We also note that only $\mathcal{C}=\omega(1)$ is interesting; otherwise, an optimal algorithm (such as sawtooth backoff [13, 10, 2]) for the standard model will be asymptotically optimal. To capture the worst-case performance, $\mathcal{C}$ is chosen by the adversary with knowledge of $n$, subject to $\mathcal{C}=O({\texttt{poly}({n})})$, where the degree is of the polynomial is unknown. Note that this reveals no information about $n$, whose value is unknown to the packets a priori.
Given any wakeup algorithm, the two pertinent costs are: (i) the total cost of collisions, which is the number of collisions incurred over the execution of the algorithm multiplied by $\mathcal{C}$, and (ii) the latency. We report on both of these metrics separately, with the aim of minimizing their maximum.
Notation. An event holds with high probability (w.h.p.) in $x$ if it occurs with probability at least $1-O(1/{\texttt{poly}({x})})$, where we can make the constant degree of the polynomial arbitrarily large. We use $\log^{c}y$ to denote $(\log y)^{c}$.
@@ 1.2 Our Results
The performance of our algorithm, Aim-High (AH), is parameterized by $\epsilon$, which is a small positive constant set by the user prior to execution.
@@
AH solves the wakeup problem with the following costs:
@@
    If $\mathcal{C}=\Omega(\log^{\Theta(1/\epsilon)}n)$, then the expected latency is $O(\mathcal{C}^{\frac{1}{2}+2\epsilon})$ and the expected total cost of collisions is $O(\mathcal{C}^{\frac{1}{2}+\epsilon})$.
@@
    Else, the expected latency is $O(\log^{4+\frac{1}{2\epsilon}}n)$ and the expected total cost of collisions is $O(\log^{4+\frac{3}{2\epsilon})}n)$.
By comparison, existing approaches to wakeup incur a collision cost that is $\Omega(\mathcal{C})$ (see Section 3.1). In contrast, Theorem 1.1 implies that if $\mathcal{C}$ is “large”, AH offers performance that is sublinear in $\mathcal{C}$. Else, latency and expected total cost of collisions are polylogarithmic-in-$n$, which echoes prior wakeup results (see Section 2).
Many upper bounds are achieved by “fair” algorithms. A fair algorithm is defined by a sequence $p_{1}$, $p_{2}$, $\dots$, where $p_{i}$ denotes the sending probability used by every packet in the $i$-th slot. In Section 5, we establish the following trade-off between latency and expected total cost of collisions.
@@
Let $\alpha$ and $\beta$ be constants such that $1/(2\epsilon)\geq\alpha\geq 0$ and $\beta\geq 0$. Any fair algorithm for the wakeup problem that has $O(\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n)$ latency has an expected total cost of collisions $\Omega(\mathcal{C}^{(1/2)-\alpha\epsilon}/\log^{\beta}n)$.
@@ 2 Related Work
Wakeup, or leader election, addresses the fundamental problem of symmetry breaking. Early work by Willard [20] establishes a $\lg\lg n+O(1)$ upper bound with a matching lower bound for fair algorithms under collision detection (CD).
@@
With multiple channels and collision detection, Fineman et al. [9] give an efficient contention-resolution algorithm. In the SINR model [17], Fineman et al. [8] achieve logarithmic latency and establish matching lower bounds. Chang et al. [6] examine energy efficiency under varying feedback models, while Gilbert et al. [11] study the effect of system-size prediction, deriving corresponding bounds.
Research on the cost of collisions is still in its early stages. Anderton et al. [1] show that collisions create substantial discrepancies between theoretical and observed performance in backoff algorithms, underscoring the need to account for collision costs. Building on this, Biswas et al. [4] provide one of the first analyses, addressing the problem of contention resolution and deriving bounds on expected latency and total expected collision cost.
@@ 4 $w_{0}\leftarrow 2^{\mathcal{C}^{\epsilon}}$
5 $j\leftarrow 1$
6
7
8// Iterations
9 while true do
10
11 // Halving Phase
@@ 12 $w_{\mbox{\tiny cur}}\leftarrow w_{0}$
14    for slot $i=1$ to $d\sqrt{\mathcal{C}}\,\ln(w_{\mbox{\tiny cur}})$ do
15
16
17     -2ptSend packet with probability $1/w_{cur}$
@@ 21       $w_{\mbox{\tiny cur}}\leftarrow w_{\mbox{\tiny cur}}/2$
22
23
24 // Doubling Phase
25 $w_{\mbox{\tiny cur}}\leftarrow w_{0}$
26 for $s=1$ to $2^{j}$ do
@@ 27    $w_{\mbox{\tiny cur}}\leftarrow 2w_{\mbox{\tiny cur}}$
28    for slot $i=1$ to $d\sqrt{\mathcal{C}}\,\ln(w_{\mbox{\tiny cur}})$ do
29
30
31     -2ptSend packet with probability $1/w_{cur}$
@@ 35 $j\leftarrow j+1$
3 Algorithm Specification
The pseudocode for AH is given above, and each packet executes this algorithm upon being activated. Under AH, each packet begins with an initial window size of $w_{0}$ $2^{\mathcal{C}^{\epsilon}}$.
[…]
@@ 3 Algorithm Specification
Doubling phase. At the beginning of this phase, $w_{\mbox{\tiny cur}}$ starts at $2w_{0}$. Each packet sends with probability $\Theta(1/w_{\mbox{\tiny cur}})$ in each slot of a sample (Line 1 and 1). If a success occurs, the execution terminates; otherwise, the window doubles and another sample is executed. This process continues until $2^{j}$ samples (each corresponding to a doubling of the window) are performed, where $j=1$ initially, and increments at the end of each iteration.
@@ 3.1 Context and Overview
We start by considering how prior approaches fare in a model when there is a collision cost. Then, we discuss our design choices regarding AH, along with an overview of our analysis.
@@
A natural approach for breaking symmetry is to “backoff”, where each packet sends with probability $1/2^{i}$ in slot $i\geq 0$. At $i=\lg(n)$, the probability of a success in a slot becomes constant. An example of its use in solving wakeup is given by Jurdziński and Stachowiak [14], who achieve $O(n/\log n)$ latency with bounded error $\varepsilon$ in a similar model to ours (without a notion of collision cost); however, each active packet starts with a high (i.e., constant) sending probability, which inflicts $\Omega(\mathcal{C})$ cost.
So, we can observe that while backoff may be an important part of any solution, a key obstacle is how to set the initial sending probability suitably low. Since $n$ is unknown a priori, it seems that no matter what initial sending probability we choose—even if it depends on the known value of $\mathcal{C}$—the adversary may choose a sufficiently large value of $n$ such that collisions are unavoidable. What can we do?
@@
Setting an Initial Window Size. The initial sending probability needs to be set carefully. Too high and we incur many collisions as we discussed above.
Consider a window of size $w^{*}=\Theta(n\sqrt{\mathcal{C}})$; we refer to this as the size of a good window. Informally, once packets reach a good window, the chance of a success in any slot of the window is at least $\binom{n}{1}(1/w^{*})\approx 1/\sqrt{\mathcal{C}}$. Therefore, over $\Theta(\sqrt{\mathcal{C}})$ slots, we expect a success. At the same time, the probability of a collision is approximately $\binom{n}{2}(1/w^{*})^{2}\approx 1/\mathcal{C}$, and thus the expected collision cost per slot is $1$, and $O(\sqrt{\mathcal{C}})$ over $\Theta(\sqrt{\mathcal{C}})$ slots. This observation dictates the size of our samples, which are $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ and appear in lines 1 and 1 in the halving phase, and lines 1 and 1 in the doubling phase of our pseudocode.
In designing AH, we hope that $w^{*}\leq w_{0}$, since in this case, the above reasoning yields good expected collision cost (and, as we will show, good latency). However, since the adversary sets $\mathcal{C}$ with knowledge of $n$, our hopes may be dashed; the adversary can always make collisions overwhelmingly likely. However, in this case, our consolation prize is that collisions are cheap, and the degree of cheapness corresponds to $w_{0}$. We choose to "aim high" with the initial window size; that is, setting $w_{0}=2^{\mathcal{C}^{\epsilon}}$. This choice sets up our analysis of the following two cases:
@@
    If $n\sqrt{\mathcal{C}}\leq w_{0}$, then the good-window size is less than or equal to the initial-window size. Here, the probability of a collision is not too high. Thus, so long as we adjust the sending probability upwards in a conservative manner, we can reduce expected collision cost; this is the purpose of the halving phase. Of course, there is a tension with latency, since this adjustment cannot be too conservative, otherwise the latency will be too large.
@@
    Else $n\sqrt{\mathcal{C}}>w_{0}$, and the good-window size is larger than the initial-window size. Here, the collision probability may be “too high”. However, a key realization is that $\mathcal{C}$ is small; we will argue that $n>2^{\mathcal{C}^{\epsilon}}/\sqrt{\mathcal{C}}$ implies that $\mathcal{C}=O(\log^{1/\epsilon}n)$. Thus, we can tolerate these collisions as we increase the window size in the doubling phase.
Why Iterations? Although AH is likely to terminate when a good window is reached, there is some small probability of failure. Suppose we are in Case 1, but this failure occurs in the halving phase. Then, this small probability of failure yields good expected total cost of collisions as the window size decreases past the good window down to $2$ (see Lemma 3). However, if the halving phase is never revisited, the packets may accrue intolerable latency by executing the doubling phase without success (as the sending probability drops off exponentially fast). Conversely, suppose we are in Case 2 and this failure occurs. Then, if we never reset $w_{\mbox{\tiny cur}}$ to (roughly) $w_{0}$, the latency costs may again be high. Therefore, iterations facilitate good expected latency (see Lemma 6).
Parameter $\epsilon$. Throughout, we treat $\epsilon$ as a small constant for the purposes of our analysis. Yet, it is worth pointing out how it interacts with the latency and the expected total cost of collisions. The former is improved by making $\epsilon$ small, since this shortens the halving phase. So, if aiming high pays off, then a smaller $\epsilon$ can reduce latency. However, we pay for this in terms of the latter metric; specifically, when $n\sqrt{\mathcal{C}}>w_{0}$ and the algorithm looks to succeed in the doubling phase, we have $\mathcal{C}=O(\log^{1/\epsilon}n)$. Thus, degree of the polynomial increases with the reciprocal of $\epsilon$.
4 Upper Bound
In this section, we analyze the performance of AH. Define $\eta$ $\max\{n,\mathcal{C}\}$. A good window was loosely discussed in Section 3.1.2; we define it as any window with size in $n\sqrt{\mathcal{C}},3n\sqrt{\mathcal{C}}]$. Throughout, we assume that $n$ is sufficiently large. We make use of the following well-known facts.